# Establishing a Foundation for Tetun Text Ad-Hoc Retrieval: Stemming, Indexing, Retrieval, and Ranking

### v2

@@ Abstract.
Searching for information on the internet and digital platforms to satisfy an information need requires effective retrieval solutions. However, such solutions are not yet available for Tetun, making it challenging to find relevant documents for text-based search queries in this language. To address these challenges, we investigate Tetun text retrieval with a focus on the ad-hoc retrieval task. The study begins by developing essential language resources—including a list of stopwords, a stemmer, and a test collection—which serve as foundational components for solutions tailored to Tetun text retrieval. Various strategies are then explored using both document titles and content to evaluate retrieval effectiveness. The results show that retrieving document titles, after removing hyphens and apostrophes without applying stemming, significantly improves retrieval performance compared to the baseline. Efficiency increases by 31.37%, while effectiveness achieves an average gain of 9.40% in MAP@10 and 30.35% in nDCG@10 with DFR BM25. Beyond the top-10 cutoff point, Hiemstra LM demonstrates strong performance across various retrieval strategies and evaluation metrics. Contributions of this work include the development of Labadain-Stopwords (a list of 160 Tetun stopwords), Labadain-Stemmer (a Tetun stemmer with three variants), and Labadain-Avaliadór (a Tetun test collection containing 59 topics, 33,550 documents, and 5,900 qrels).
@@ 1. Introduction
Ad-hoc text retrieval is the task of retrieving documents from large text collections in response to user queries, without prior knowledge of the topics that users are likely to search, highlighting the unpredictable nature and short duration of each search (Voorhees and Harman, 1998; Voorhees, 2004). Users typically express their information needs through natural language text queries and submit them to a search system. The retrieval system then retrieves, ranks, and returns documents relevant to the query, presenting the most relevant documents at the top of the list, with less relevant ones further down.
@@
These challenges are also evident in the development of resources for Tetun, a LRL spoken by over 923,000 people in Timor-Leste (de Jesus, 2023). Timor-Leste is a Southeast Asian island country characterized by its multilingualism, comprises two official languages (Tetun and Portuguese), two working languages (English and Indonesian) (Vasconcelos et al., 2011), and over 30 dialects spoken across the territory (de Jesus, 2023). Tetun which was a dialect, became one of Timor-Leste’s official languages when the country restored its independence in 2002 (Vasconcelos et al., 2011). Despite its status as an official language, Tetun is characterized by data scarcity, with fewer than 45,000 documents available on the web as of 2023 (Kudugunta et al., 2023; de Jesus and Nunes, 2024a). Moreover, Tetun is a less studied and computerized language, lacking essential resources for effective text retrieval, including a stopword list, a stemmer, and a test collection for the ad-hoc retrieval task.
To tackle the aforementioned challenges, we investigate strategies for Tetun text ad-hoc retrieval, including evaluating the impact of stemming and stopwords, to develop the most effective retrieval solutions for Tetun. The research questions (RQs) we addressed in this study are the following:
@@
RQ2
How can query and document preprocessing improve text retrieval effectiveness in Tetun?
Given that Tetun words contain accented letters (á, é, í, ó, ú, ñ), apostrophes (‘), and hyphens for mono-semantic compound words, our objective is to investigate the impact of query and document preprocessing on the effectiveness of text retrieval in Tetun text-based search. In response to the research questions above, we hypothesize that applying language-specific preprocessing components to queries and documents can improve retrieval effectiveness without the need for stemming, particularly in retrieving short texts such as document titles. This hypothesis is grounded in findings from our preliminary study on Tetun text ad-hoc retrieval (de Jesus, 2022).
To test this hypothesis, we began by developing a list of Tetun stopwords (Labadain-Stopwords), a Tetun stemmer (Labadain-Stemmer), and a Tetun test collection (Labadain-Avaliadór), using the Labadain-30k+ dataset (de Jesus and Nunes, 2024a). For Labadain-Stemmer, three variants were developed: light, moderate, and heavy. The Labadain-Stemmer performance was evaluated both as standalone systems and for their impact within the retrieval system, known as intrinsic and extrinsic assessments (Mollá and Hutchinson, 2003; Jones and Galliers, 1996). The Labadain-Avaliadór was developed following TREC guidelines and assessed by native Tetun-speaking students. This collection was then used to evaluate the retrieval effectiveness for Tetun text ad-hoc retrieval. The evaluation results reveal that query and document preprocessing, specifically the removal of hyphens and apostrophes, improves overall retrieval effectiveness, without the need for stemming, thus confirming our hypothesis. Contributions of this work include (i) the development of Labadain-Stopwords, (ii) Labadain-Stemmer with three variants, (iii) Labadain-Avaliadór, and (iv) baselines for Tetun text ad-hoc retrieval.
The remaining sections of this paper are organized as follows. Section 2 provides background and related work. An overview of Tetun and its linguistic characteristics is outlined in Section 3. Section 4 describes the dataset used in this study, while Section 5 outlines the phases of Tetun text ad-hoc retrieval development. Section 6 details the process of constructing Labadain-Stopwords, and Section 7 describes the development of the Labadain-Stemmer. The construction of Labadain-Avaliadór is explained in Section 8. Section 9 presents the baselines for Tetun text ad-hoc retrieval. Finally, Section 10 summarizes the conclusions and discusses potential future work.
@@ 2. Background and Related Work
Text retrieval process includes several preprocessing steps, such as removing stopwords and stemming. The key topics in this study include the development of stopwords, stemming, test collection for evaluation, and baselines for Tetun text ad-hoc retrieval. We present an overview and related work for each of these topics in the following subsections.
@@ 2.1. Stopwords
Fox (Fox, 1990) applied a term frequency approach to the Brown Corpus to generate a stopword list for English, a method that has since been widely adopted to develop lists for other languages, including French (Savoy, 1999); Marathi, Czech, Hungarian, and other five languages (Ghosh and Bhattacharya, 2017); and Kinyarwanda and Kirundi (Niyongabo et al., 2020). Furthermore, Lo et al. (Lo et al., 2005) introduced the normalized inverse document frequency (NIDF) for stopword detection in experiments with four English TREC collections, finding NIDF more effective than TF and IDF. Later, Ferilli (Ferilli, 2021) proposed term-document frequency (TDF) and, after testing it with two Italian corpora, reported TDF as more effective than TF, IDF, and NIDF, particularly for smaller datasets. More recently, Ali et al. (Ali et al., 2024) demonstrated that the network-based approach leveraging topological properties of co-occurrence networks, such as in-degree, out-degree, and degree, outperforms traditional term-weighting techniques, with in-degree providing the most consistent results in both high- and low-resource languages, including Tetun.
In information retrieval, stopwords generally contribute minimal value to retrieving relevant documents for a given query (Baeza-Yates and Ribeiro-Neto, 2011; Manning et al., 2009). Therefore, removing these words from both queries and documents can enhance retrieval efficiency and effectiveness (Sahu et al., 2023; Sahu and Pal, 2023; Baeza-Yates and Ribeiro-Neto, 2011; Croft et al., 2009). Despite this, other studies have reported that the effectiveness of stopword removal varies between ranking models and languages, demonstrating that it is beneficial in some cases but not in others (Ghosh and Bhattacharya, 2017; Dolamic and Savoy, 2010; Savoy, 1999). In a study with French, Savoy (Savoy, 1999) found that retaining stopwords performed better than removing them when using the BM25 ranking model. Similarly, Dolamic and Savoy (Dolamic and Savoy, 2010) did not observe significant differences in retrieval effectiveness for Marathi and Bengali, while for Hindi, including stopwords improved retrieval effectiveness, with average gains of approximately 20%. Ghosh and Bhattacharya (Ghosh and Bhattacharya, 2017) further demonstrated variability in retrieval effectiveness across datasets within the same language, observing that stopword removal did not lead to noticeable differences in retrieval effectiveness when tested with datasets from FIRE, CLEF, and TREC collections across several languages.
@@ 2.2. Stemming
Stemming is an essential component of text processing that captures the relationship between different variations and forms of a word resulting from inflection (e.g., plurals, tenses, and gender) or derivation (e.g., converting a verb into a noun by adding suffixes) and reduces them to a common root (Croft et al., 2009). A stem is the root form of a word that remains after the removal of its affixes (Baeza-Yates and Ribeiro-Neto, 2011). For example, the words [ car, cars, car’s, cars’ ] are all reduced to “car”. Stemming can be useful for improving retrieval effectiveness by minimizing index size and reducing the number of distinct terms.
@@
The first stemming algorithm was proposed by Lovins (Lovins, 1968), based on the principles of iteration and longest match. The iteration principle assumes that affixes are attached to stems in a particular order from a predefined set of affixes. The algorithm removes affixes from either the beginning or the end of the word, depending on which affix class is detected. According to the longest match principle, if multiple endings within a class match, the longest one should be removed. Since then, various stemming techniques have been developed, including rule-based, dictionary-based, and automatic stemmers. One of the most notable examples is the Porter Stemmer (Porter, 1980), a widely used rule-based suffix removal algorithm for English stemming due to its simplicity and performance (Baeza-Yates and Ribeiro-Neto, 2011).
@@
Several Asian languages have adopted suffix-stripping stemmers based on the Porter and Lovins approaches, including Sanskrit (Sahu and Pal, 2023), Sundanese (Ardiyanti et al., 2018), Czech (Dolamic and Savoy, 2009), and Indonesian (Adriani et al., 2007). An advanced version of the Porter Stemmer is the Snowball stemming algorithm, which supports multiple languages, including Portuguese, Spanish, and German (Snowball, 2002). Snowball applies a set of predefined stemming rules tailored to the specific morphological structure of each language, primarily focusing on removing suffixes to reduce words to their root forms.
@@
In an experiment on French, Savoy (Savoy, 1999) found stemming particularly beneficial for retrieving short documents (scientific abstracts averaging 24.5 indexing terms), with improvements in average precision (AP) in various retrieval models. For longer documents (news articles averaging 182.2 terms per article) or cases where accents were ignored, stemming yielded only marginal benefits. Similarly, Braschler and Ripplinger (Braschler and Ripplinger, 2004), using German data from CLEF 2000 and 2001 datasets, reported stemming gains in mean average precision (MAP) of up to 23% for short queries (title only) and up to 11% for long queries (a combination of title, description, and narrative).
@@
However, Hollink et al. (Hollink et al., 2004), in a monolingual document retrieval experiment across eight languages using the CLEF 2002 dataset, found that stemming improved MAP for Finnish, French, German, and Swedish but did not improve the retrieval effectiveness for Dutch, English, Italian, or Spanish. Likewise, Flores and Moreira (Flores and Moreira, 2016), in an experiment with four different languages from the CLEF 2005 and 2006 datasets, reported that stemming was generally beneficial in MAP for Portuguese, French, and Spanish but not for English when tested with different stemmers.
@@
In studies focused on Asian LRLs, Sahu and Pal (Sahu and Pal, 2023), in an experiment with Sanskrit, reported that stemming improved retrieval effectiveness by an overall 4.31% in MAP across six different ranking models. Sahu et al. (Sahu et al., 2023) found a 1.41% improvement in the MAP score in an experiment with Urdu when tested with various stemming approaches. Similarly, Adriani et al. (Adriani et al., 2007) reported a 2. 00% improvement in the MAP score in an experiment with Indonesian.
@@ 2.3. Test Collection
Relevance judgments typically fall into two categories: binary and graded relevance. Binary relevance categorizes each document as either relevant or non-relevant to the user’s query, assigning a score of 1 for relevant and 0 for non-relevant documents. The graded relevance evaluates documents on multiple levels of relevance, with the most relevant documents awarding higher scores. Binary relevance is predominantly used for experimental research in the TREC collections. In the TREC-9 Web Track, three-level graded relevance was introduced: not relevant, relevant, and highly relevant (Hawking, 2000). Later, Sormunen proposed a four-level relevance scale consisting of non-relevant, marginally relevant, relevant, and highly relevant (Sormunen, 2002). Kekäläinen (Kekäläinen, 2005) adopted a similar four-point scale, but labeled the third level as “fairly relevant” rather than “relevant”. This four-point scale was subsequently implemented across multiple TREC tracks. The ad-hoc retrieval task was a central focus of the TREC tracks held from 1992 to 1999 and was revisited on the robust track from 2003 to 2005 (Voorhees, 2006).
The TREC-style approach, derived from the Cranfield paradigms, is commonly used for developing test collections for LRLs. Sahu and Pal (Sahu and Pal, 2023) employed this method to create a Sanskrit test collection, comprising 7,057 news articles and 50 topics, with query formulation and relevance judgments conducted by two Ph.D. students. Similarly, Chavula and Suleman (Chavula and Suleman, 2021) developed a test collection for three Bantu languages—Chichewa, Citumbuka, and Cinyanja—utilizing documents sourced from newspapers, Wikipedia, and web pages. Their collection contains 13,627 documents and 387 topics, with query formulation and relevance assessment carried out by six recruited assessors. Moreover, Aleahmad et al. (AleAhmad et al., 2009) developed a Hamshahri test collection for Persian, using the news corpus consisting of 166,774 documents, 65 queries. The queries formulation and relevance assessment were carried out by 17 university students.

### v6

# Establishing a Foundation for Tetun Ad-hoc Text Retrieval: Stemming, Indexing, Retrieval, and Ranking

@@ Abstract.
Searching for information on the internet and digital platforms requires effective retrieval solutions. However, such solutions are not yet available for Tetun, making it difficult to find relevant documents for search queries in this language. To address this gap, we investigate Tetun text retrieval with a focus on the ad-hoc retrieval task. The study begins with the development of essential language resources—including a list of stopwords, a stemmer, and a test collection—that serve as a foundation for Tetun text retrieval. Various strategies are evaluated using document titles and content. The results show that retrieving document titles, after removing hyphens and apostrophes but without applying stemming, improves performance compared to the baseline. Efficiency increases by 31.37%, while effectiveness achieves an average relative gains of +9.40% in MAP@10 and +30.35% in NDCG@10 with DFR BM25. Beyond the top-10 cutoff point, Hiemstra LM demonstrates strong performance across multiple retrieval strategies and evaluation metrics. The contributions of this work include the development of Labadain-Stopwords (a list of 160 Tetun stopwords), Labadain-Stemmer (a Tetun stemmer with three variants), and Labadain-Avaliadór (a Tetun test collection comprising 59 topics, 33,550 documents, and 5,900 qrels). These resources are publicly available to support future research in Tetun information retrieval.
@@ 1. Introduction
Ad-hoc text retrieval is the task of retrieving documents from large text collections in response to user queries without prior knowledge of the topics that users are likely to search, highlighting the unpredictable nature and short duration of each search (Voorhees and Harman, 1998; Voorhees, 2004). Users typically express their information needs through natural language text queries and submit them to a search system. The retrieval system then retrieves, ranks, and returns documents relevant to the query, presenting the most relevant documents at the top of the list, with less relevant ones further down.
@@
These challenges are also faced in the development of resources for Tetun, a LRL spoken by over 923,000 people in Timor-Leste (de Jesus, 2023). Timor-Leste is a Southeast Asian island country characterized by its multilingualism, comprising two official languages (Tetun and Portuguese), two working languages (English and Indonesian) (Vasconcelos et al., 2011), and over 30 dialects spoken across the territory (de Jesus, 2023). Tetun, which was a dialect, became one of Timor-Leste’s official languages when the country restored its independence in 2002 (Vasconcelos et al., 2011). Despite its status as an official language, Tetun is characterized by data scarcity, with fewer than 45,000 documents available on the web as of 2023 (Kudugunta et al., 2023; de Jesus and Nunes, 2024a). Moreover, Tetun is a less-studied and computerized language, lacking essential resources for effective text retrieval, including a stopword list, a stemmer, and a test collection for the ad-hoc retrieval task.
To tackle the aforementioned challenges, we investigated strategies for Tetun ad-hoc text retrieval, including evaluating the impact of stemming and stopwords, to identify the most effective retrieval solutions for Tetun. The research questions (RQs) we addressed in this study are the following:
@@
How can text preprocessing techniques tailored to Tetun’s linguistic characteristics improve retrieval effectiveness?
RQ2
@@
Given that Tetun words contain accented letters (á, é, í, ó, ú, ñ), apostrophes (‘), and hyphens in monosemantic compound words, our objective is to investigate the impact of query and document preprocessing on the effectiveness of text retrieval in Tetun text-based search. In line with the research questions above, we hypothesize that applying language-specific preprocessing to queries and documents can improve retrieval effectiveness without the need for stemming, particularly when retrieving short texts such as document titles. This hypothesis is grounded in findings from our preliminary study on Tetun ad-hoc text retrieval, which reported a 3.1% relative improvement in overall MAP when stemming was not applied (de Jesus, 2022).
To test this hypothesis, we began by developing a list of Tetun stopwords (Labadain-Stopwords), a Tetun stemmer (Labadain-Stemmer), and a Tetun test collection (Labadain-Avaliadór) using the Labadain-30k+ dataset (de Jesus and Nunes, 2024a). For Labadain-Stemmer, three variants were developed: light, moderate, and heavy. The Labadain-Stemmer performance was evaluated both as standalone systems and for their impact within the retrieval system (intrinsic) and extrinsic assessments (Mollá and Hutchinson, 2003; Jones and Galliers, 1996). The Labadain-Avaliadór was developed following TREC guidelines and assessed by native Tetun-speaking students. This collection was then used to evaluate the retrieval effectiveness for Tetun ad-hoc text retrieval. The contributions of this work include: (i) the development of Labadain-Stopwords, (ii) the creation of Labadain-Stemmer with three variants, (iii) the construction of Labadain-Avaliadór, and (iv) the establishment of baselines for Tetun ad-hoc text retrieval.
The remainder of this paper is organized as follows. Section 2 reviews background and related work. Section 3 provides an overview of Tetun and its linguistic characteristics. The dataset used in this study is described in Section 4, and Section 5 outlines the methodology for establishing baselines in Tetun ad-hoc text retrieval. Section 6 presents the construction of Labadain-Stopwords, while Section 7 describes the development of Labadain-Stemmer. Section 8 details the creation of Labadain-Avaliadór, and Section 9 reports the baselines for Tetun ad-hoc text retrieval. Finally, Section 10 concludes the paper and discusses directions for future work.
@@ 2. Background and Related Work
Text retrieval typically involves several preprocessing steps, such as stopword removal and stemming. The key topics relevant to this study are the development of stopword lists, stemming approaches, test collections for evaluation, and baselines for Tetun ad-hoc text retrieval. Background on each of these topics is provided in the following subsections.
@@ 2.1. Stopwords
Fox (1990) applied a term frequency approach to the Brown Corpus to generate a stopword list for English, a method that has since been widely adopted to develop stopwords for other languages, including French (Savoy, 1999); Marathi, Czech, Hungarian, and five other languages (Ghosh and Bhattacharya, 2017); and Kinyarwanda and Kirundi (Niyongabo et al., 2020).
Furthermore, Lo et al. (2005) introduced the normalized inverse document frequency (NIDF) for stopword detection, evaluating it on four English TREC collections and showing that NIDF outperforms TF and IDF. Later, Ferilli (Ferilli, 2021) proposed the term-document frequency (TDF) metric and, after testing it on two Italian corpora, found that TDF surpasses TF, IDF, and NIDF, particularly in smaller datasets. More recently, Ali et al. (2024) demonstrated that a network-based approach exploiting topological properties of co-occurrence networks—such as in-degree, out-degree, and degree—outperforms traditional term-weighting techniques, with in-degree yielding the most consistent results across both high- and low-resource languages, including Tetun.
In IR, stopwords generally contribute minimal value to retrieving relevant documents for a given query (Baeza-Yates and Ribeiro-Neto, 2011; Manning et al., 2009). Therefore, removing these words from both queries and documents can enhance retrieval efficiency and effectiveness (Sahu et al., 2023; Sahu and Pal, 2023; Baeza-Yates and Ribeiro-Neto, 2011; Croft et al., 2009). Despite this, other studies have reported that the effectiveness of stopword removal varies between ranking models and languages, demonstrating that it is beneficial in some cases but not in others (Ghosh and Bhattacharya, 2017; Dolamic and Savoy, 2010; Savoy, 1999). In a study on French, Savoy (1999) found that retaining stopwords performed better than removing them when using BM25. Similarly, Dolamic and Savoy (2010) did not observe significant differences in retrieval effectiveness for Marathi and Bengali, while for Hindi, including stopwords improved retrieval effectiveness, with average gains of approximately 20% in mean average precision (MAP). Ghosh and Bhattacharya (2017) further demonstrated variability in retrieval effectiveness across datasets within the same language, observing that stopword removal did not lead to noticeable differences in retrieval effectiveness when evaluated on datasets from FIRE, CLEF, and TREC collections across several languages.
@@ 2.2. Stemming
Stemming is an essential component of text processing that captures the relationship between different variations and forms of a word resulting from inflection (e.g., plurals, tenses, and gender) or derivation (e.g., converting a verb into a noun by adding suffixes) and reduces them to a common root (Croft et al., 2009). A stem is the root form of a word that remains after the removal of its affixes (Baeza-Yates and Ribeiro-Neto, 2011). Stemming can be useful for improving retrieval effectiveness by minimizing index size and reducing the number of distinct terms.
@@
The first stemming algorithm was proposed by Lovins (1968), based on the principles of iteration and longest match. The iteration principle assumes that affixes are attached to stems in a particular order from a predefined set of affixes. The algorithm removes affixes from either the beginning or the end of the word, depending on which affix class is detected. According to the longest match principle, if multiple endings within a class match, the longest one should be removed. Since then, various stemming techniques have been developed, including rule-based, dictionary-based, and automatic stemmers. One of the most notable examples is the Porter Stemmer (Porter, 1980), a widely used rule-based suffix removal algorithm for English stemming due to its simplicity and performance (Baeza-Yates and Ribeiro-Neto, 2011).
@@
Several Asian languages have adopted suffix-stripping stemmers based on the Porter and Lovins approaches, including Sanskrit (Sahu and Pal, 2023), Sundanese (Ardiyanti et al., 2018), Czech (Dolamic and Savoy, 2009), and Indonesian (Adriani et al., 2007). An advanced version of the Porter Stemmer is the Snowball stemming algorithm, which supports multiple languages, including Portuguese, Spanish, and German (Snowball, 2002). Snowball applies a set of predefined stemming rules tailored to the specific morphological structure of each language, primarily focusing on removing suffixes.
@@
In an experiment on French, Savoy (1999) found that stemming was particularly beneficial for retrieving short documents (scientific abstracts averaging 24.5 indexing terms), with improvements in average precision (AP) in various retrieval models. For longer documents (news articles averaging 182.2 terms per article) or cases where accents were ignored, stemming yielded only marginal benefits. Similarly, Braschler and Ripplinger (2004), evaluating on German data from the CLEF 2000 and 2001 datasets, reported stemming gains in MAP of up to 23% for short queries (title only) and up to 11% for long queries (a combination of title, description, and narrative).
@@
However, Hollink et al. (2004), in a monolingual document retrieval experiment using the CLEF 2002 dataset across eight languages, reported inconsistent results. Specifically, stemming improved MAP for Finnish, French, German, and Swedish but had no positive effect on Dutch, English, Italian, or Spanish. Likewise, Flores and Moreira (2016), in an experiment with four different languages from the CLEF 2005 and 2006 datasets, reported that stemming was generally beneficial in MAP for Portuguese, French, and Spanish but not for English when tested with different stemmers.
@@
In studies on Asian LRLs, Sahu and Pal (2023) reported that stemming improved retrieval effectiveness for Sanskrit by 4.31% in MAP across multiple ranking models. Likewise, Sahu et al. (2023) observed a 1.41% MAP improvement for Urdu when testing several stemming approaches, while Adriani et al. (2007) found a 2.00% MAP improvement in experiments with Indonesian.
@@ 2.3. Test Collection
Relevance judgments typically fall into two categories: binary and graded relevance. Binary relevance categorizes each document as either relevant or non-relevant to the user’s query, assigning a score of 1 for relevant and 0 for non-relevant documents. The graded relevance evaluates documents on multiple levels of relevance, with the most relevant documents awarding higher scores. Binary relevance is predominantly used for experimental research in the TREC collections. In the TREC-9 Web Track, three-level graded relevance was introduced: not relevant, relevant, and highly relevant (Hawking, 2000). Later, Sormunen proposed a four-level relevance scale consisting of non-relevant, marginally relevant, relevant, and highly relevant (Sormunen, 2002). Kekäläinen (2005) adopted a similar four-point scale but labeled the third level as “fairly relevant” rather than “relevant”. This four-point scale was subsequently implemented across multiple TREC tracks. The ad-hoc retrieval task was a central focus of the TREC tracks held from 1992 to 1999 and was revisited on the robust track from 2003 to 2005 (Voorhees, 2006).
The TREC-style approach, derived from the Cranfield paradigm, is commonly used to develop test collections for LRLs. Sahu and Pal (2023) applied this method to create a Sanskrit test collection comprising 7,057 news articles and 50 topics, with queries and relevance judgments produced by two Ph.D. students. Similarly, Chavula and Suleman (2021) constructed a test collection for three Bantu languages—Chichewa, Citumbuka, and Cinyanja—using documents from newspapers, Wikipedia, and web pages. Their collection includes 13,627 documents and 387 topics, with queries and relevance assessments carried out by six recruited assessors. Furthermore, AleAhmad et al. (2009) developed the Hamshahri test collection for Persian, based on a news corpus of 166,774 documents and 65 queries, with queries and relevance assessments performed by 17 volunteer students.